OpenStudy (anonymous):

Which of the following expressions is equivalent to 6m - 10? 2(3m + 5) -2(5 - 3m) 2(3m - 10)

jhonyy9 (jhonyy9):

so check this 6m = 3*2m 10 = 2*5 what can you factorizing out ?

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

B. -2(5 - 3m)

OpenStudy (anonymous):

u shure

OpenStudy (anonymous):

ofc

OpenStudy (anonymous):

http://www.wolframalpha.com/ <----- i go their when i have problems with math

jhonyy9 (jhonyy9):

yes try calculi it -2*5=-10 and -2*(-3)=6

OpenStudy (anonymous):

i got your answer from their

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=6m+-+10

OpenStudy (anonymous):

take look

OpenStudy (anonymous):

its a then i enterd it and it came out with a

OpenStudy (anonymous):

it also shows work

OpenStudy (mathstudent55):

Take each choice and multiply it out using the distributive property. A. 2(3m + 5) = 6m + 10 (not correct) B. -2(5 - 3m) = -10 + 6m = 6m - 10 (looking good) C. 2(3m - 10) = 6m - 20 (not good) B. is the correct choice.

jhonyy9 (jhonyy9):

with a you not can getting minus

OpenStudy (anonymous):

on the site for me it said the awnser was a

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=6m+-+10 <------ look here all stuff explains perfect when i had problems that web helped me

OpenStudy (anonymous):

you prob put numbers wrong or asked wrong

OpenStudy (anonymous):

any ways link above is with your problem

OpenStudy (anonymous):

6m - 10

OpenStudy (anonymous):

thats waht i put in

OpenStudy (anonymous):

take look on screen

OpenStudy (anonymous):

alternate form